I am trying to solve p2 – 15p = –54
but I am not getting the answers provided.
I used the quadratic equation
-b +- (sqrt) b^2 - 4ac / 2a
-(-15) +- (sqrt) (-15)^2 - 4(1)(54)
15 +- (sqrt) 225-216 / 2
15 +- 9 / 2
X = 12 X = 3
what did I do wrong?

Answer by stanbon(75887)

solve p2 – 15p = –54
p^2 - 15p + 54 = 0
(P-7)(p-8) = 0
p = 7 or p = 8
==================
Cheers,
Stan H.